OpenStudy (anonymous):

If x+2+y what is the value of lx-yl+ly-xl

OpenStudy (anonymous):

If x + 2 + y? wtf does that even mean.

OpenStudy (anonymous):

I guess he means x + 2 + y = 0, in which case, your answer = 4 + 4y

OpenStudy (anonymous):

oops sorry its suppose to be x+2=y

OpenStudy (anonymous):

Just plut it in to the formula. So, if y = x + 2, then x - y = x - x - 2 = -2 and y - x = 2, so your answer is 4 (absolute value of -2 is 2)

OpenStudy (anonymous):

That makes more sense.For the record, Max, if the x / y didn't cancel in the | |, then you couldn't say what it was; it would depend on whether x/x >= 0

OpenStudy (anonymous):

x and y*

OpenStudy (anonymous):

That is correct Newton

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

oops sorry its suppose to be x+2=y